# BOSC: A Backdoor-based Framework for Open Set Synthetic Image Attribution

### v1

@@ IV-D Evaluation metrics
The performance in a closed-set scenario is evaluated by providing the Accuracy (Acc) when the model is tested on in-set samples. Open-set performance is evaluated in terms of rejection accuracy (out-of-set detection performance) and classification performance. Specifically, the capability to distinguish between in-set and out-of-set samples is assessed by computing the ROC curve and measuring the area under the curve (AU-ROC).
In addition, the capability to retain and correctly classify in-set samples is also measured. Following prior works in this field [26, 15, 37], we consider the Open-Set Classification Rate (OSCR) curve, and measure the area under this curve (AU-OSCR). Formally, let $D_{k}$ indicate the set of in-set samples, and $D_{u}$ the set of out-of-set test samples. Let the event that a sample comes from an in-set (out-of-set) class be the negative (positive) event. The True Positive Rate (TPR) and False Positive Rate (FPR) for a given rejection threshold $\nu$, are defined as
@@
The ROC curve plots the TPR vs FPR values while the OSCR curve plots CCR vs FPR, obtained by varying the threshold $\nu$. By adding the constraint on the correct classification of known samples, the OSCR metric measures the trade-off between open and closed-set performance and provides an overall evaluation of the entire system.
@@ V Experimental Results
TABLE II: Performance of architecture attribution in closed-set (Accuracy) and open-set (AU-ROC, AU-OSCR) settings. The best results are shown in bold (the second-best is underlined).
@@
Accuracy AU-ROC AU-OSCR Accuracy AU-ROC AU-OSCR Accuracy AU-ROC AU-OSCR Accuracy AU-ROC AU-OSCR
ARPL [13] 100 79.59 79.55 99.94 77.84 77.78 99.98 83.01 83.00 99.97 80.15 80.11
AKPF [15] 99.95 75.27 75.27 100 88.98 88.98 99.56 91.89 91.6 99.84 85.38 85.28
PCSSR [14] 99.47 83.11 82.88 99.57 68.58 68.50 98.55 71.32 70.74 99.20 74.34 74.04
RCSSR [14] 99.62 82.65 82.46 99.21 57.98 57.84 98.65 70.79 70.32 99.16 70.47 70.21
Res50Vit [30] 99 79 78.32 99 76 75.89 99 68 67.93 99 74.33 74.05
SiaVerify [40] 100 82.24 82.31 100 82.44 82.41 100 82.98 82.89 100 82.55 82.54
POSE [37] 98.56 75.97 75.60 96.90 86.73 85.53 96.70 83.00 81.50 97.39 81.90 80.88
Baseline 100 82.62 82.56 100 73.10 73.10 99.99 65.99 65.98 99.99 73.90 73.88
BOSC (Ours) 100 95.31 95.31 99.95 95.43 95.41 99.96 90.00 89.99 99.96 93.58 93.42
Figure 6: Robustness to image processing. From left to right: JPEG compression, brightness, contrast and saturation change, and Gaussian blur. From top to bottom: Closed-set: Accuracy, Open-set: AU-ROC and AU-OSCR. Only JPEG compression was considered for augmentation during training.
@@ V-A Performance analysis
The closed-set and open-set performance of the proposed method in all three configurations is reported in Table II, together with the performance of the state-of-the-art methods described in section IV-C. In addition, to better assess the gain achieved by BOSC, we also report the performance achieved using the same EfficientNet-B4 baseline to build an $N$-class classifier and adopting the MLS for the detection of out-of-set samples (as mentioned in Section II-A, MLS has been proven to achieve the best rejection performance in many cases [16], and is adopted for open-set classification in several papers [16, 30]). The table referred to this method as ’baseline’. We see that all methods achieve nearly perfect accuracy in the closed-set scenario, while the open-set performance are different. In particular, CSSR methods show limited effectiveness, as well as ResVit50. The best performing state-of-the-art method is AKPF, which achieves an average AU-ROC equal to 85.38% and an AU-OSCR equal to 85.28%. Most of the methods, especially the general OSR methods, exhibit unstable open-set performance across the three configurations, e.g. for AKPF, the AU-ROC ranges from 75.27% in Config S1 to 91.89% in Config S3. Results are more stable for SiaVerify and POSE.
Regarding BOSC, it achieves the best open-set performance (AU-ROC = 93.58% and AU-OSCR= 93.42% on average), with very limited variability across the configurations. In particular, BOSC outperforms the best-performing state-of-the-art method AKPF with a gain of 8.20% and 8.14% in AU-ROC and AU-OSCR, respectively. The baseline is also significantly outperformed by BOSC, with a gain of approximately 20% in both AU-ROC and AU-OSCR, confirming the effectiveness of the backdoor-based framework we have introduced.
@@ V-B Robustness to image processing manipulations
Fig. 6 reports the closed-set and open-set performance. BOSC has robustness similar to state-of-the-art methods for in-set samples, while it gets superior performance in the open-set case, with a gain in AU-ROC and AU-OSCR larger than 10% in all the cases. This confirms the intuition that, since the trigger images are not affected by the processing (being superimposed at test time), their features are affected to a lesser extent by processing. The most critical case is JPEG compression, notwithstanding its inclusion in the training set.
V-C Architecture vs model attribution
@@
We also evaluated the capability of BOSC of correctly attributing to the source architecture images generated by unknown models, that is, models different than those considered during training yet sharing the same in-set architectures. In particular, the generative models used for these tests are obtained from the in-set architectures considering: i) different training strategies (for StyleGAN2, LSGM and DDPM) ii) different training datasets of real images (for Latent diffusion and Taming transformer), and iii) different image resolutions (for StyleGAN3). The details of the mismatch between the models used for training and testing are provided in Table III. With regard to the models obtained considering a mismatch in the training strategy, StyleGAN-ada refers to training with adaptive augmentation for the discriminator [51], while DDPM-ema is obtained by training with the exponential moving average strategy [52]. Finally, for LSGM, the number of stages for the training is changed, from the default value 2 to 3. In the third stage, re-training is performed by training only the SGM prior, leaving the Nouveau VAE (NVAE) component fixed [^6] See https://github.com/NVlabs/LSGM for the details.. The models obtained with the 2-stage and 3-stage training are referred to as quantitative and qualitative models, respectively.
@@
TABLE III: Closed-set and Open-set results in the case of models trained with different datasets, parameters, and training procedures (in the last line, the number in the models’ names indicates the image resolution). The underlined performance refers to the average results across different configurations.
Architecture Type of Mismatch Train Test ACC AU-OSCR
DDPM (Config-S1) Training Methodology DDPM DDPM-ema 100 85.89
StyleGAN2 (Config-S1&S2&S3) Training Methodology StyleGAN2-f StyleGAN-ada 99.84 92.15
Taming Transformer (Config-S1) Real Dataset CelebA FFHQ 98.20 82.53
Latent Diffusion (Config-S1&S2) Real Dataset CelebA FFHQ 77.63 71.3
LSGM (Config-S2) Training Methodology Quantitative (2-stages) Qualitative (3-stages) 99.60 76.46
StyleGAN3 (Config-S3) Image Resolution StyleGAN3 t-1024& t-ffhqu1024&r StyleGAN3 t-ffhqu256 99.35 78.17
@@
The last two columns of Table III report the closed-set accuracy and the AU-OSCR obtained when the system is tested with the out-of-set models. The results show that the closed-set performance is very good (Acc above 98%) in all cases, with the exception of Latent Diffusion, where we get Acc = 77.63%. The model mismatch affects the open-set performance more, and in fact, the AU-OSCR computed on the mismatched samples decreases. Performance is great in the case of Syle-GAN2 and remains pretty good also for DDPM and Taming Transformer, while they drop in the case of Latent Diffusion, LSGM and SyleGAN3. A strategy that we expect can help mitigating this issue is to include multiple models for every architecture inside the training set. Arguably, doing so should induce the system to better learn the model variability for a given generative architecture.
@@ V-D1 Choice of rejection score
TABLE IV: Ablation study on the effect of mixup augmentation.
@@
BOSC (w/- Mixup) 99.97 93.58 93.57
@@
TABLE V: Summary of the 19 editing classes (18 + ’None’) and the configurations of in-set (11 classes) and out-of-set (8 classes).
@@
StyleCLIP Expression (T5, T6): Angry, Surprised Hair style (T7-T12): Afro, Purple_hair, Curly_hair, Mohawk, Bobcut, Bowlcut Identity change (T13-T18): Taylor_swift, Beyonce, Hilary_clinton, Trump, Zuckerberg, Depp
@@
Config F1 In-set: T4, T1, T2, T5, T6, T7, T8, T9, T10, T11, T12 Out-of-set: T0, T3, T13, T14, T15, T16, T17, T18
Config F2 T5, T1, T2, T3, T4, T11, T12, T13, T14, T15, T18 Out-of-set: T0, T6, T7, T8, T9, T10, T16, T17
Config F3 T2, T1, T3, T4, T6, T10, T12, T15, T16, T17, T18 Out-of-set: T0, T5, T7, T8, T9, T11, T13, T14
@@ V-D2 Mixup augmentation
We also ran some experiments to assess the benefit of the mixup augmentation strategy. Table IV reports the closed-set and open-set performance achieved by BOSC when training is carried out with and without mixup augmentation. The performance of the baseline is also reported. We see that, while the accuracy values of all the models are the same, the open-set performance improve significantly with the adoption of mixup augmentation. In particular, the gain brought by the mixup strategy in the performance of BOSC is 7.58% in both AU-ROC and AU-OSCR.
@@ VI Application to facial editing classification
In this section, we describe the experiments we ran on an additional image forensic task, namely the open-set classification of facial attribute editing [7].
The classification of facial attribute editing consists in detecting the particular facial attribute modified by means of a conditional generative network. Possibly manipulate facial attributes, including age, gender, or expression. Classifying the manipulated attributes in facial images enables effective analysis and detection of tampered or manipulated content, ensuring the integrity and authenticity of visual data in digital environments.

### v2

@@ IV-D Evaluation metrics
The performance in a closed-set scenario is evaluated by providing the Accuracy (Acc) when the model is tested on samples from in-set clasases. Open-set performance is evaluated in terms of rejection accuracy (out-of-set detection performance) and classification performance. Specifically, the capability to distinguish between in-set and out-of-set classes is assessed by computing the ROC curve and measuring the area under the curve (AU-ROC).
In addition, the capability to retain and correctly classify in-set class samples is also measured. Following prior works in this field [31, 18, 42], we consider the Open-Set Classification Rate (OSCR) curve, and measure the area under this curve (AU-OSCR). Formally, let $D_{k}$ indicate the set of in-set samples, and $D_{u}$ the set of out-of-set test samples. Let the event that a sample comes from an in-set (out-of-set) class be the negative (positive) event. The True Positive Rate (TPR) and False Positive Rate (FPR) for a given rejection threshold $\nu$, are defined as
@@
The ROC curve plots the TPR vs FPR values while the OSCR curve plots CCR vs FNR (where FNR = 100% - TPR), obtained by varying the threshold $\nu$. [^7] We point that, in the literature, the OSCR is often defined as the CCR vs FPR, because the event that a sample comes from an in-set class is taken as the positive event. By adding the constraint on the correct classification of known samples, the OSCR metric measures the trade-off between open and closed-set performance and provides an overall evaluation of the entire system. For an overall characterization of the open-set classification performance of the system, we also report the TPR and CCR values obtained by fixing the threshold $\nu$ based on the FPR.
@@ V Experimental Results
TABLE II: Performance of architecture attribution in closed-set (Accuracy) and open-set (AU-ROC, AU-OSCR) settings. The best results are shown in bold (the second-best is underlined). The average EER of out-of-set detection is also reported.
@@
Accuracy AU-ROC AU-OSCR Accuracy AU-ROC AU-OSCR Accuracy AU-ROC AU-OSCR Accuracy AU-ROC EER AU-OSCR
ARPL [16] 100 79.59 79.55 99.94 77.84 77.78 99.98 83.01 83.00 99.97 80.15 33.81 80.11
AKPF [18] 99.95 75.27 75.27 100 88.98 88.98 99.56 91.89 91.6 99.84 85.38 24.76 85.28
PCSSR [17] 99.47 83.11 82.88 99.57 68.58 68.50 98.55 71.32 70.74 99.20 74.34 34.11 74.04
RCSSR [17] 99.62 82.65 82.46 99.21 57.98 57.84 98.65 70.79 70.32 99.16 70.47 25.69 70.21
Res50Vit [35] 99 79 78.32 99 76 75.89 99 68 67.93 99 74.33 32.50 74.05
SiaVerify [45] 100 82.24 82.31 100 82.44 82.41 100 82.98 82.89 100 82.55 27.47 82.54
POSE [42] 98.56 75.97 75.60 96.90 86.73 85.53 96.70 83.00 81.50 97.39 81.90 26.30 80.88
Baseline 100 82.62 82.56 100 73.10 73.10 99.99 65.99 65.98 99.99 73.90 33.64 73.88
BOSC (Ours) 100 95.31 95.31 99.95 95.43 95.41 99.96 90.00 89.99 99.96 93.58 14.65 93.42
Figure 6: Robustness to image processing. From left to right: JPEG compression, brightness, contrast and saturation change, and Gaussian blur. From top to bottom: Closed-set: Accuracy, Open-set: AU-ROC and AU-OSCR. The metric in the Y-axis on the left applies to every plot in the row. A vertical dashed line separates JPEG compression, which was considered as augmentation method during training, from the other processing, not considered during training.
TABLE III: TPR and CCR when FPR=5% for architecture attribution in open-set. The values of the threshold $\nu$ are also reported. The best results are shown in bold (the second-best is underlined).
Methods Config S1 Config S2 Config S3 Average
Open-set Open-set Open-set Open-set
TPR CCR $\nu$ TPR CCR $\nu$ TPR CCR $\nu$ TPR CCR
ARPL [16] 37.0 94.9 15.2 41.2 94.9 15.0 63.1 94.8 14.8 47.1 94.8
AKPF [18] 25.2 95.4 -17.7 58.4 95.5 9.4 68.5 95.2 11.5 50.7 95.4
PCSSR [17] 31.7 91.3 -5.2 17.1 92.0 1.5 12.9 86.9 2.7 20.6 90.1
RCSSR [17] 22.2 90.7 -5.1 7.9 92.3 -5.7 12.4 87.9 -3.6 14.1 90.3
Res50Vit [35] 38.5 94.4 8.9 42.6 94.6 10.5 54.2 94.7 8.9 45.1 94.6
SiaVerify [45] 45.4 94.4 0.5 45.3 94.8 0.5 45.4 94.4 0.5 45.3 94.5
POSE [42] 32.3 94.6 0.5 42.3 93.5 0.4 49.8 93.5 0.5 41.5 93.9
Baseline 65.4 95.5 11.8 35.5 95.5 7.7 56.7 95.5 11.7 52.5 95.5
BOSC (Ours) 79.6 95.5 13.2 80.2 95.5 11.3 59.6 95.5 12.0 73.2 95.5
@@ V-A Performance analysis
The closed-set and open-set performance of the proposed method in all three configurations is reported in Table II, together with the performance of the state-of-the-art methods described in Section IV-C. We see that all methods achieve nearly perfect accuracy in the closed-set scenario, while the open-set performance are different. In particular, CSSR methods show limited effectiveness, as well as ResVit50. The best open-set results are achieved with AKPF, SiaVerify and POSE, with the last two being specifically developed for OSIA. Notably, the best performing state-of-the-art method is AKPF, which achieves an average AU-ROC equal to 85.38% and an AU-OSCR equal to 85.28%. However, this method exhibit very unstable open-set performance across the three configurations, e.g. for AKPF, the AU-ROC ranges from 75.27% in Config S1 to 91.89% in Config S3, while results tend to be more stable for SiaVerify and POSE. A possible reason is that, being based on the synthesis of open-set samples, AKPF is more sensitive to the composition of the unknown class and performance are very good when in-set and out-of-set classes are very different. In Config S3, where AKPF got the best performance, only GAN architectures are included in the in-set, while LD and transformers architectures are all in the out-of-set, thus corresponding to an easy configuration Regarding BOSC, it achieves the best open-set performance (AU-ROC = 93.58% and AU-OSCR= 93.42% on average), with very limited variability across the configurations. In particular, BOSC outperforms the best-performing state-of-the-art method AKPF with a gain of 8.20% and 8.14% in AU-ROC and AU-OSCR, respectively. The baseline is also significantly outperformed, with a gain of approximately 20% in both AU-ROC and AU-OSCR, confirming the effectiveness of the backdoor-based framework we have introduced. The table also reports the average Equal Error Rate (EER) of the ROC curve in the second-last column, which in our case is 10% lower than AKPF (second best method). The TPR and CCR achieved by the various methods when the threshold $\nu$ is fixed to achieve the target FPR= 5% are reported in Table III, along with the corresponding values of $\nu$. We see that BOSC achieves both the highest TPR and CCR. The TPR in particular is more than 20% larger than for the other methods.
Fig. 7 reports the confusion matrices of open-set classification for the BOSC method for all the configurations, showing the distribution of the errors among the classes, when FPR=5%. The diagonal (1-5) reports the CCR for every in-set class, while the last diagonal element shows the classification accuracy (accuracy of rejection) of the samples from out-of-set, i.e. the TPR. We observe that, for samples from in-set classes, the classification errors are always in favor of the rejection class, i.e., all the samples which are accepted as in-set are then correctly classified.
Overall, the results we got show that BOSC significantly improves the performance of OSIA. Although it is difficult to understand exactly where the advantage of BOSC comes from, the main reasons for its good performance can be identified as follows. With respect to methods based on output score thresholding (like MSP and MLS thresholding), the creation of a link between in-set classes and triggers allows to differentiate better between samples from out-of-set classes and hard-to-classify samples of in-set classes, to which the closed-set model assigns a low confidence. Another advantage comes from the reduction of the open space risk. The necessity of recognizing the presence of the triggers, whose content departs significantly from the content of the query images, forces the classifier to enlarge the latent space, hence increasing its capacity to distinguish between samples form in-set and out-of-set classes. Eventually, we argue that the necessity to create a distinct link between the samples of different classes and the various triggers, forces the classifier to compact the representation of in-set classes in the latent space, hence contributing to further reduce the open space risk.
Figure 7: Confusion matrices of the BOSC method for open-set classification, when FPR is fixed to 5%.
@@ V-B Robustness to image processing manipulations
Fig. 6 reports the closed-set and open-set performance. BOSC has robustness similar to state-of-the-art methods in the closed-set scenario, while it gets superior performance in the open-set case, with a gain in AU-ROC and AU-OSCR larger than 10% in all the cases. This confirms the intuition that, since the trigger images are not affected by the processing (being superimposed at test time), their features are affected to a lesser extent by processing. The most critical case is JPEG compression, notwithstanding its inclusion in the training set.
V-C Generalization analysis
@@
We also evaluated the capability of BOSC of correctly attributing to the source architecture images generated by unknown models, that is, models different than those considered during training yet sharing the same in-set architectures In particular, the generative models used for these tests are obtained from the in-set architectures considering: i) different training strategies (for StyleGAN2, LSGM and DDPM), ii) different training datasets of real images (for Latent diffusion and Taming transformer), and iii) different image resolutions (for StyleGAN3). The details of the mismatch between the generative models considered to get the images used for training and testing are provided in Table IV (column 2, 3, and 4). With regard to the models obtained considering a mismatch in the training strategy, StyleGAN-ada refers to training with adaptive augmentation for the discriminator [56], while DDPM-ema is obtained by training with the exponential moving average strategy [57]. Finally, for LSGM, the number of stages for the training is changed, from the default value 2 to 3. In the third stage, re-training is performed by training only the SGM prior, leaving the Nouveau VAE (NVAE) component fixed [^8] See https://github.com/NVlabs/LSGM for the details.. The models obtained with the 2-stage and 3-stage training are referred to as quantitative and qualitative models, respectively.
@@
TABLE IV: Results (closed and open-set) in the presence of a mismatch between the generative models considered during training and testing for a same architecture. Generative models considered for testing are trained considering different datasets, parameters, and training procedures w.r.t. the models considered during training (in the last line, the number in the models’ names indicates the image resolution). The underlined performance refers to the average results across different configurations.
Architecture
Type of
Mismatch
Train Test ACC AU-OSCR
DDPM
(Config-S1)
Training
Methodology
DDPM DDPM-ema 100 85.89
StyleGAN2
(Config-S1&S2&S3)
Training
Methodology
StyleGAN2-f StyleGAN-ada 99.84 92.15
Taming Transformer
(Config-S1)
Real Dataset CelebA FFHQ 98.20 82.53
Latent Diffusion
(Config-S1&S2)
Real Dataset CelebA FFHQ 77.63 71.3
LSGM
(Config-S2)
Training
Methodology
Quantitative
(2-stages)
Qualitative
(3-stages)
99.60 76.46
StyleGAN3
(Config-S3)
Image Resolution
StyleGAN3
t-1024&
t-ffhqu1024&r
StyleGAN3
t-ffhqu256
99.35 78.17
@@
The last two columns of Table IV report the closed-set accuracy and the AU-OSCR obtained for the specific in-set architecture (detailed in column 1), considering test images obtained from mismatched generative models. The results show that the closed-set performance is very good (Acc above 98%) in all cases, with the exception of Latent Diffusion, where we get Acc = 77.63%. This confirms that the classifier indeed looks at architecture-level features, and not at some model specific features. The model mismatch affects the open-set performance more, and in fact, the AU-OSCR computed on the mismatched samples decreases. Performance is great in the case of StyleGAN2 and remains pretty good also for DDPM and Taming Transformer, while they drop in the case of Latent Diffusion, LSGM and StyleGAN3. A strategy that we expect can help mitigating this issue is to include multiple models for every architecture inside the training set. Arguably, doing so should induce the system to better learn the model variability for a given generative architecture.
Finally, we also ran an experiment to measure the performance in a realistic scenario where the OSIA methods are tested with pristine images. Specifically, in this experiment, we included images from two pristine datasets, namely the CelebA and the FFHQ (2.000 images per each), in the out-of-set class, which then comprises a mixture of pristine images and images generated by different generative architectures (2 pristine datasets and 5 generative architectures). Fig. 9 shows the results, averaged across the three configurations, in terms of AU-ROC, AU-OSCR, EER and TPR at FPR = 5% (the FP threshold is set on the in-set class, hence it is the same as before). We see that the inclusion of pristine images leads to a performance drop for many methods, especially the best performing ones. Notably, BOSC remains the method achieving the best results according to all metrics.
Figure 9: Average open-set performance (in terms of AU-ROC, AU-OSCR, TPR@FPR5% and EER) of the various methods when pristine images (from CelebaHQ and FFHQ) are also included in the out-of-set class. The results without including pristine images are also reported.
@@ V-D1 Choice of rejection score
TABLE V: Ablation study on the effect of mixup augmentation.
@@
BOSC (w/- Mixup) 99.97 93.58 93.42
Figure 11: A different set of trigger images used in the ablation study for the attribution task.
@@
TABLE VI: Summary of the 19 editing classes (18 + ’None’) and the configurations of in-set (11 classes) and out-of-set (8 classes).
@@
StyleCLIP
Expression (T5, T6): Angry, Surprised
Hair style (T7-T12): Afro, Purple_hair,
Curly_hair, Mohawk, Bobcut, Bowlcut
Identity change (T13-T18): Taylor_swift, Beyonce,
Hilary_clinton, Trump, Zuckerberg, Depp
@@
Config F1
In-set: T4, T1, T2, T5, T6, T7, T8, T9, T10, T11, T12
Out-of-set: T0, T3, T13, T14, T15, T16, T17, T18
Config F2
T5, T1, T2, T3, T4, T11, T12, T13, T14, T15, T18
Out-of-set: T0, T6, T7, T8, T9, T10, T16, T17
Config F3
T2, T1, T3, T4, T6, T10, T12, T15, T16, T17, T18
Out-of-set: T0, T5, T7, T8, T9, T11, T13, T14
@@ V-D2 Mixup augmentation
We also ran some experiments to assess the benefit of the mixup augmentation strategy. Table V reports the closed-set and open-set performance achieved by BOSC when training is carried out with and without mixup augmentation. The performance of the baseline is also reported. We see that, while the accuracy values of all the models are the same, the open-set performance improve significantly with the adoption of mixup augmentation. In particular, the gain brought by the mixup strategy in the performance of BOSC is 7.58% in both AU-ROC and AU-OSCR.
V-D3 Choice of Triggers
Finally, we run additional experiments with different trigger images, namely, using geometrical trigger patterns (shown in Fig. 11). We got the following average performance: Accuracy 99.62%, AU-ROC = 77.47%, AU-OSCR = 77.30%. It is worth noticing that, although this set of trigger images do not perform so well as the cartoon images (see Table V), with a reduction of approx 15% in both AU-ROC and AU-OSCR), the open-set performance remains superior to those achieved by the baseline, which is surpassed by 3/4% in both AU-ROC and AU-OSCR.
@@ VI Application to facial editing classification
In this section, we describe the experiments we ran and the results we got on an different image forensic task, namely the open-set classification of facial attribute editing [10], to validate the generality of the BOSC framework. The classification of facial attribute editing consists in detecting the particular facial attribute modified by means of a conditional generative network. Manipulated facial attributes include age, gender, expression or hair color. Classifying the manipulated attributes in facial images enables effective analysis and detection of tampered or manipulated content, ensuring the integrity and authenticity of visual data in digital environments.